# Fortune: Formula-Driven Reinforcement Learning for Symbolic Table Reasoning in Language Models

### v2

@@ Abstract
Tables are a fundamental structure for organizing and analyzing data, making effective table understanding a critical capability for intelligent systems. While large language models (LMs) demonstrate strong general reasoning abilities, they continue to struggle with accurate numerical or symbolic reasoning over tabular data, especially in complex scenarios. Spreadsheet formulas provide a powerful and expressive medium for representing executable symbolic operations, encoding rich reasoning patterns that remain largely underutilized. In this paper, we propose Formula Tuning (Fortune), a reinforcement learning (RL) framework that trains LMs to generate executable spreadsheet formulas for question answering over general tabular data. Formula Tuning reduces the reliance on supervised formula annotations by using binary answer correctness as a reward signal, guiding the model to learn formula derivation through reasoning. We provide a theoretical analysis of its advantages and demonstrate its effectiveness through extensive experiments on seven table reasoning benchmarks. Formula Tuning substantially enhances LM performance, particularly on multi-step numerical and symbolic reasoning tasks, enabling a 7B model to outperform OpenAI o1 on table understanding. This highlights the potential of formula-driven RL to advance symbolic table reasoning in LMs.   Corresponding authors.
@@
1 Introduction
Tables are a common and practical data structure in daily life, playing a central role in data collection, representation, and analysis [1, 2]. Recent advances in large language models (LLMs) [3, 4, 5] have brought impressive performance across a wide range of natural language processing tasks, including language understanding [6, 7] and general reasoning [8]. Naturally, LLMs have also been applied to tabular data understanding and reasoning [9, 10, 11].
However, reasoning over tabular data remains a key challenge for language models (LMs) [11], primarily because of the numerical nature and intricate structure of tables, which pose significant obstacles to understanding both content and layout. Besides, some findings [12] suggest current LLMs rely on pattern memorization over genuine rule learning, often leading to incorrect mathematical computations during traditional chain-of-thought reasoning, also referred to as textual reasoning. Some approaches introduce symbolic methods [13, 14], where models first generate symbolic representations (such as programs) and then execute them to obtain results. While this improves arithmetic accuracy, such methods often struggle to generalize due to limited symbolic reasoning or program generation capabilities. Instead of truly understanding the context and generating problem-solving programs, LMs often fall back on memorized code snippets from pretraining [15]. Since not all complex symbolic patterns can be memorized, and given that high-quality code supervision is scarce [16], more effective strategies for symbolic table reasoning are needed.
Recent progress in reinforcement learning (RL) for LMs shows promise for overcoming such limitations. For example, DeepSeek-R1 [17] improves mathematical reasoning in LLMs via rule-based RL rewards, without relying on step-by-step annotations. DeepRetrieval [18] uses retrieval metrics as RL rewards to train models to reason over queries that maximize real-world retrieval performance across search engines and databases. DeepCoder [19] also demonstrates the effectiveness of RL for code reasoning and generation. These works collectively demonstrate the promise of RL in enabling LMs to perform robust symbolic reasoning without explicit intermediate supervision.
When it comes to symbolic table reasoning, spreadsheet formula [20] is a powerful and versatile tool. In real-world scenarios, tabular data is often stored in spreadsheet formats (e.g., Microsoft Excel, Google Sheets) [21], where each cell holds individual data values. The spreadsheet formulas embedded in these files act as lightweight, program-like constructs that enable users to compute, transform, and reason over data. Compared to structured interfaces such as SQL or Python/Pandas, which are typically restricted to relational or flat table schemas, spreadsheet formulas offer greater flexibility, as they can be applied to arbitrary two-dimensional tables without structural constraints [22]. Moreover, spreadsheet formulas are Turing complete [23], making them particularly well-suited as a medium for symbolic table reasoning.
We envision that by training LMs to understand and generate spreadsheet formulas, they can acquire more robust and generalizable symbolic reasoning capabilities over tabular data. However, current LLMs still struggle to produce accurate and reliable spreadsheet formulas, as highlighted by recent evaluations [24]. At the same time, existing publicly available spreadsheet datasets tend to include relatively simple formulas [25], rely on heuristic conversions from SQL [26], or synthesize formulas using LLMs in constrained question-answering settings [22]. These approaches fall short of capturing the complexity necessary for diverse symbolic reasoning tasks and real-world downstream applications. This limitation poses a significant barrier to effectively leveraging spreadsheet formulas for training LMs in symbolic table reasoning.
To address these challenges, we propose Formula Tuning (Fortune), a RL framework designed to teach LMs to perform symbolic reasoning over general tabular data through spreadsheet formula. Specifically, our framework leverages answer correctness of formula execution results as a reward signal to guide the LMs in deriving formulas through reasoning (Figure 1). This approach reduces reliance on supervised formula annotations and enables LMs to generate executable formulas that answer questions over tables with improved accuracy. Extensive experiments validate the effectiveness of Formula Tuning, demonstrating that RL is more effective than supervised fine-tuning (SFT) in enhancing the symbolic reasoning capabilities of LMs. We also find that initializing RL with SFT as a cold start [17] provides a stronger foundation and raises the upper bound of RL performance, with SFT serving as a form of knowledge injection. Notably, this enables a 7B model to outperform OpenAI o1 in overall performance (68.48% vs. 66.90%). Furthermore, we train both textual and symbolic reasoning components using RL in Fortune++. By jointly leveraging both components during inference, our method achieves strong performance across multiple benchmarks (e.g., 82.54% on WikiTQ, 95.06% on TabFact, 87.24% on HiTab, 80.47% on FinQA, and 93.20% on AIT-QA).
In summary, this paper makes the following key contributions:
  • We propose Formula Tuning (Fortune), a reinforcement learning framework that enhances symbolic reasoning for table understanding by training language models to generate executable spreadsheet formulas.
  • We provide a theoretical analysis and discussion comparing textual versus symbolic reasoning in table understanding, as well as supervised fine-tuning versus reinforcement learning in symbolic table reasoning.
  • We conduct extensive experiments on seven table understanding benchmarks, demonstrating the effectiveness of Formula Tuning, and perform comprehensive analyses to provide deeper insights.
2 Related Work
Table Understanding and Reasoning. Many studies have explored fine-tuning language models (LMs) to improve their ability to understand and reason over tabular data. Building on the masked language modeling introduced by BERT [27], models such as TaPas [28], PaSTA [29], and TUTA [30] propose specialized pretraining strategies tailored for tables. TAPEX [31] pretrains an encoder-decoder model as a neural SQL executor to better capture the semantics of table operations. Recent efforts, including TableLLaMA [32] and TableGPT [33], build upon large decoder-only language models pretrained for general-purpose table understanding across a wide range of downstream tasks.
Other studies focus on enabling LMs to better perform table-related tasks without fine-tuning. For example, Dater [34] proposes strategies for dynamically constructing sub-tables, modifying the input context to enhance comprehension. Chain-of-Table [35] models table reasoning as a sequence of transformations using predefined operations, gradually generating sub-tables to support complex multi-step inference. TableMaster [11] introduces a general framework for table understanding and underscores the importance of symbolic reasoning in handling complex scenarios. Given the structured and often numerical nature of tabular data, program-of-thought prompting [13] and other symbolic approaches [36, 37, 38] have demonstrated strong effectiveness for table reasoning.
Formula Learning. A growing body of research has explored the potential of spreadsheet formulas as a powerful means to enhance table understanding. NL2Formula [26] constructs a formula generation dataset by converting text-to-SQL tasks into spreadsheet formulas, enabling position-aware reasoning from natural language queries. ForTap [39] leverages spreadsheet formulas as pretraining signals to enhance numerical reasoning. Auto-Formula [40] applies contrastive learning to transfer formulas from similar spreadsheets for formula recommendation. SpreadsheetCoder [41] formulates formula prediction as a program synthesis task, leveraging both headers and surrounding cell values for context. FLAME [42] trains a small domain-specific model tailored for formula repair and completion. TabAF [22] jointly generates answers and formulas for table question answering, but relies on supervised fine-tuning over datasets generated by LLMs.
Reinforcement Learning for Language Models. Reinforcement Learning (RL) [43] is a machine learning paradigm that trains agents to make decisions through interaction with an environment, with the goal of maximizing cumulative rewards. In the era of large language models (LLMs), RL has gained significant traction as an effective framework for aligning models with human preferences. A prominent example is Reinforcement Learning from Human Feedback (RLHF) [44, 45, 46], which leverages the Proximal Policy Optimization (PPO) algorithm [47] and human preference data to train a reward model that guides the fine-tuning of LLMs. Building on RLHF, more recent algorithms such as GRPO [48] and REINFORCE++ [49] aim to enhance reward modeling and mitigate issues like biased optimization [50].
Reasoning with Language Models. It has been observed that sufficiently large language models (LMs) can demonstrate emergent reasoning capabilities [51, 52]. Chain-of-thought prompting [53] is one technique used to elicit step-by-step reasoning, significantly improving performance on complex tasks. Further advances include self-consistency [54] and structuring the reasoning process in forms like trees [55] or graphs [56, 57] are also useful for more complex reasoning tasks. RL has also been used to directly improve reasoning skills during training [58, 59]. Notably, DeepSeek-R1 [17] demonstrates that large-scale RL can substantially boost the reasoning abilities of LMs. In terms of application, DeepRetrieval [18] applies RL to teach models how to reason about interacting with search engines for information retrieval, while DeepCoder [19] uses RL for code reasoning and generation tasks. Rec-R1 [60] also bridges LLMs and recommendation systems through RL.
Figure 2: Comparison of Textual versus Symbolic Reasoning in Table Understanding, and Supervised Fine-Tuning (SFT) versus Reinforcement Learning (RL) in Symbolic Table Reasoning.
@@ 3 Methodology
In this section, we present a theoretical analysis and discussion comparing textual versus symbolic reasoning in table understanding, as well as supervised fine-tuning (SFT) versus reinforcement learning (RL) in symbolic table reasoning (Figure 2). We then introduce our proposed training framework, Formula Tuning. All notations are list at Appendix M.
@@ 3.1 Task Formulation
where each $C_{i,j}$ may contain a data value, structural information (e.g., a top header or a left header), or be empty. In practice, we linearize $\mathbb{T}$ into a text sequence before feeding it to the LM.
The LM then generates an output $a\in\mathcal{A}$, which can be either a final textual answer or a spreadsheet formula $f$ that produces the answer upon execution. Our goal is to optimize the parameters $\theta$ that maximize the expected table-understanding performance, measured by a reward function $r(a\mid s)$. Formally,
@@
3.2 Textual vs. Symbolic Reasoning in Table Understanding
Definition 1 (Textual and Symbolic Policies).
  1. 1.
  2. 2.
Lemma 1 (Reward Decomposition).
For textual reasoning, the expected reward is:
$\mathbb{E}_{f\sim\pi_{\theta}^{\mathrm{sym}}}[r(\operatorname{exec}(f,\mathbb{T})\mid s)]=\sum_{f}\pi_{\theta}^{\mathrm{sym}}(f\mid s)\cdot\mathbbm{1}[\operatorname{exec}(f,\mathbb{T})=a^{\star}(s)].$ (4)
Assumption 1 (Symbolic Reasoning Setting).
  1. 1.
  2. 2.
  3. 3.
    The textual and symbolic policies are equally capable of planning high-level solution strategies in their respective formats (text or formula).
Theorem 1 (Symbolic Reasoning Superiority).
Under Assumption 1, the expected reward achieved by symbolic reasoning is greater than or equal to that of textual reasoning for any input $s$:
The proof of Theorem 1 is provided in Appendix C.1.
Remark 1 (Symbolic Reasoning Advantage).
Maximizing the expected reward in Eq. (2) therefore favors the symbolic policy $\pi_{\theta}^{\mathrm{sym}}$ over the textual policy $\pi_{\theta}^{\mathrm{txt}}$. When both policies have equal capacity for high-level solution planning, symbolic reasoning offers distinct advantages: it is better suited for handling complex tables and questions that require intricate reasoning and intensive calculations. As a result, it can achieve higher accuracy than models that rely solely on free-text numerical manipulation.
3.3 SFT vs. RL in Symbolic Table Reasoning
Assumption 2 (SFT Setting).
  1. 1.
    Sufficient expressivity. The model class $\{\pi_{\theta}(\cdot\mid s)\}$ is expressive enough to represent the teacher policy $\pi_{g}(\cdot\mid s)$, in the sense that
  2. 2.
  3. 3.
Lemma 2 (MLE Minimizes KL Divergence).
$\mathbb{E}_{a\sim\pi_{g}}\left[-\log\pi_{\theta}(a\mid s)\right]=H\left(\pi_{g}(\cdot\mid s)\right)+D_{\mathrm{KL}}\left(\pi_{g}\parallel\pi_{\theta}\right),$ (7)
The proof of Lemma 2 is provided in Appendix C.2.
Lemma 3 (Convergence of SFT and Reward Upper Bound).
Under Assumption 2, the optimal supervised fine-tuning (SFT) policy
satisfies
$\mathbb{E}_{s\sim p,\,f\sim\pi_{\theta^{\star}}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right]\leq\mathbb{E}_{s\sim p,\,f\sim\pi_{g}}\left[r(\operatorname{exec}(f,\mathbb{T})\mid s)\right].$ (10)
The proof of Lemma 3 is provided in Appendix C.3.
Remark 2 (SFT Bound).
Assumption 3 (RL Exploration).
For every input $s\in\mathcal{S}$, there exists at least one high-reward action $a^{\star}$ such that $\pi_{\theta}(a^{\star}\mid s)>0$ during training, i.e., the RL policy has a non-zero probability of exploring a correct solution.
Theorem 2 (RL Superiority).
Under Assumption 3, and assuming the reward function $r(a\mid s)$ is properly aligned with task success (e.g., exact match), reinforcement learning can achieve strictly higher expected reward than supervised fine-tuning:
$\mathbb{E}_{s\sim p,\,a\sim\pi_{\theta}^{\mathrm{RL}}}[r(a\mid s)]>\mathbb{E}_{s\sim p,\,a\sim\pi_{\theta^{\star}}^{\mathrm{SFT}}}[r(a\mid s)],$ (11)
provided that the teacher policy $\pi_{g}$ does not fully cover all correct solutions.
The proof of Theorem 2 is provided in Appendix C.4.
Remark 3 (RL Objective and Advantage).
Unlike SFT, which is constrained to imitating the teacher policy $\pi_{g}$, reinforcement learning (RL) directly maximizes the expected task reward:
This objective allows the model to discover high-reward actions that may lie outside the support of $\pi_{g}$—for example, alternative formulas that yield the correct answer but differ syntactically or structurally from those seen during supervised training.
In symbolic table reasoning, this flexibility is especially valuable. Since many distinct formulas may produce the same correct result, RL can exploit this many-to-one mapping by exploring diverse yet semantically valid expressions. As a result, RL has the potential to exceed the SFT reward bound established in Lemma 3, provided the reward signal is well aligned with the task objective and sufficient exploration occurs.

### v3

# Formula-R1: Incentivizing LLM Reasoning over Complex Tables with Numerical Computation via Formula-Driven Reinforcement Learning

@@ Abstract
Tables are a fundamental medium for organizing and analyzing data, making table reasoning a critical capability for intelligent systems. Although large language models (LLMs) exhibit strong general reasoning abilities, they still struggle with accurate numerical reasoning over tabular data, particularly in complex table settings beyond simple relational lookup. Spreadsheet formulas provide a powerful and expressive interface for executable symbolic operations, enabling rich reasoning patterns that remain largely underexplored by existing LLMs. In this paper, we introduce Formula-R1, a model trained via Formula Tuning (Fortune), a formula-driven reinforcement learning (RL) framework for table reasoning. Formula Tuning trains LLMs to generate executable spreadsheet formulas for question answering over general tabular data, using execution success and answer correctness as reward signals, thereby reducing reliance on supervised formula annotations. We demonstrate the effectiveness of Formula Tuning through extensive experiments on seven table reasoning benchmarks. It substantially improves LLM performance on table reasoning, particularly for tasks involving complex tables and multi-step numerical computation. Moreover, Formula-R1 consistently outperforms prior methods under controlled comparison settings. Beyond empirical gains, our extensive analyses provide insights into the role of RL in formula-driven table reasoning, highlighting the broader potential of formula-driven RL to enhance reasoning capabilities in LLMs.
1 Introduction
Tables are a common and practical data structure in daily life, playing a central role in data collection, representation, and analysis (He et al., 2023; Yi et al., 2025). Recent advances in large language models (LLMs) (Gunasekar et al., 2023; OpenAI, 2024; Touvron et al., 2023) have brought impressive performance across a wide range of natural language processing tasks, including language understanding (Minaee et al., 2024; Zhu et al., 2024) and general reasoning (Plaat et al., 2024). Naturally, LLMs have also been applied to tabular data understanding and reasoning (Fang et al., 2024; Zhang et al., 2024b; Cao and Liu, 2025).
However, reasoning over tabular data remains a key challenge for LLMs, primarily because of the numerical nature and intricate structure of tables, which pose significant obstacles to understanding both content and layout (Cao and Liu, 2025). Besides, some findings (Yan et al., 2025) suggest current LLMs rely on pattern memorization over genuine rule learning, often leading to incorrect mathematical computations during traditional chain-of-thought reasoning, also referred to as textual reasoning. Some approaches incorporate symbolic reasoning methods (Chen et al., 2023; Gao et al., 2023), where LLMs first generate symbolic representations, such as programs used as symbolic tools, and then execute them to obtain final results. While this improves arithmetic accuracy, such methods often struggle to generalize due to limited symbolic reasoning or program generation capabilities. Instead of truly understanding the context and generating problem-solving programs, LLMs often fall back on memorized code snippets from pretraining (Yang et al., 2024). Since not all complex symbolic patterns can be memorized, and given that high-quality code supervision is scarce (Bi et al., 2023), more effective strategies for symbolic table reasoning are needed.
Recent progress in reinforcement learning (RL) for LLMs shows promise for overcoming such limitations. For example, DeepSeek-R1 (DeepSeek-AI et al., 2025) improves mathematical reasoning in LLMs via rule-based RL rewards, without relying on step-by-step annotations. DeepRetrieval (Jiang et al., 2025) uses retrieval metrics as RL rewards to train models to reason over queries that maximize real-world retrieval performance across search engines and databases. DeepCoder (Luo et al., 2025) also demonstrates the effectiveness of RL for code reasoning and generation. These works collectively demonstrate the promise of RL in enabling LMs to perform robust symbolic reasoning without explicit intermediate supervision.
When it comes to symbolic reasoning over tables, spreadsheet formulas (Microsoft Corporation, 2025) is a powerful and versatile tool. In real-world scenarios, tabular data is often stored in spreadsheet formats (e.g., Microsoft Excel, Google Sheets) (Dong et al., 2024), where each cell holds individual data values. The spreadsheet formulas embedded in these files act as lightweight, program-like constructs that enable users to compute, transform, and reason over data. Compared to structured interfaces such as SQL or Python/Pandas, which are typically restricted to relational or flat table schemas, spreadsheet formulas offer greater flexibility, as they can be applied to arbitrary two-dimensional tables without structural constraints (Wang et al., 2025). Moreover, spreadsheet formulas are Turing complete (Smalley, 2023), making them particularly well-suited as a medium for symbolic table reasoning.
We envision that by training LLMs to understand and generate spreadsheet formulas, they can acquire more robust and generalizable symbolic reasoning capabilities over tabular data. However, current LLMs still struggle to produce accurate and reliable spreadsheet formulas, as highlighted by recent evaluations (Thorne, 2023). At the same time, existing publicly available spreadsheet datasets tend to include relatively simple formulas (Cheng et al., 2021), rely on heuristic conversions from SQL (Zhao et al., 2024a), or synthesize formulas using LLMs in constrained question-answering settings (Wang et al., 2025). These approaches fall short of capturing the complexity necessary for diverse symbolic reasoning tasks and real-world downstream applications. This limitation poses a significant barrier to effectively leveraging spreadsheet formulas for training LLMs in symbolic table reasoning.
To address these challenges, we propose Formula Tuning (Fortune), an RL framework designed to teach LLMs to perform symbolic reasoning over general tabular data through spreadsheet formulas. Specifically, our framework uses execution success and answer correctness from formula execution as reward signals to guide LLMs in deriving formulas through reasoning. This design reduces reliance on supervised formula annotations and enables LLMs to generate executable formulas that answer questions over tables more accurately. Extensive experiments on seven benchmark validate the effectiveness of Formula Tuning, showing that RL is more effective than supervised fine-tuning (SFT) at enhancing the symbolic reasoning capabilities of LLMs. We also find that initializing RL with an SFT cold start (DeepSeek-AI et al., 2025) provides a stronger foundation and improves the upper bound of RL performance, with SFT serving as a form of knowledge injection. Based on this framework, we train Formula-R1. Furthermore, we train both textual and symbolic reasoning components with RL in Formula-R1-Plus. By jointly leveraging both components during inference, our method achieves strong performance across multiple benchmarks, including 82.54% on WikiTQ, 95.06% on TabFact, 87.24% on HiTab, 80.47% on FinQA, and 93.20% on AIT-QA, outperforming previous methods on all five publicly available benchmarks.
In summary, this paper makes the following key contributions:
  • We propose Formula Tuning (Fortune), a reinforcement learning framework that enhances symbolic reasoning for table understanding by training large language models to generate executable spreadsheet formulas.
  • We conduct extensive experiments on seven table understanding benchmarks to validate the effectiveness of Formula Tuning, and we develop Formula-R1 based on this training framework.
  • We provide comprehensive empirical and theoretical analyses, offering deeper insights into textual versus symbolic reasoning for table understanding, as well as supervised fine-tuning versus reinforcement learning for symbolic table reasoning.
2 Related Work
Table Understanding and Reasoning. Many studies have explored fine-tuning language models (LMs) to improve their ability to understand and reason over tabular data. Building on the masked language modeling introduced by BERT (Devlin et al., 2019), models such as TaPas (Herzig et al., 2020), PaSTA (Gu et al., 2022), and TUTA (Wang et al., 2021) propose specialized pretraining strategies tailored for tables. TAPEX (Liu et al., 2022) pretrains an encoder-decoder model as a neural SQL executor to better capture the semantics of table operations. Recent efforts, including TableLLaMA (Zhang et al., 2024a) and TableGPT (Zha et al., 2023), build upon large decoder-only language models pretrained for general-purpose table understanding across a wide range of downstream tasks.
Other studies focus on enabling LMs to better perform table-related tasks without fine-tuning. For example, Dater (Ye et al., 2023) proposes strategies for dynamically constructing sub-tables, modifying the input context to enhance comprehension. Chain-of-Table (Wang et al., 2024) models table reasoning as a sequence of transformations using predefined operations, gradually generating sub-tables to support complex multi-step inference. TableMaster (Cao and Liu, 2025) introduces a general framework for table understanding and underscores the importance of symbolic reasoning in handling complex scenarios. Given the structured and often numerical nature of tabular data, program-of-thought prompting (Chen et al., 2023) and other symbolic approaches (Cheng et al., 2023; Nahid and Rafiei, 2024; Mao et al., 2024) have demonstrated strong effectiveness for table reasoning.
Formula Learning. A growing body of research has explored the potential of spreadsheet formulas as a powerful means to enhance table understanding. NL2Formula (Zhao et al., 2024a) constructs a formula generation dataset by converting text-to-SQL tasks into spreadsheet formulas, enabling position-aware reasoning from natural language queries. ForTap (Cheng et al., 2022) leverages spreadsheet formulas as pretraining signals to enhance numerical reasoning. Auto-Formula (Chen et al., 2024) applies contrastive learning to transfer formulas from similar spreadsheets for formula recommendation. SpreadsheetCoder (Chen et al., 2021a) formulates formula prediction as a program synthesis task, leveraging both headers and surrounding cell values for context. FLAME (Joshi et al., 2023) trains a small domain-specific model tailored for formula repair and completion. TabAF (Wang et al., 2025) jointly generates answers and formulas for table question answering, but relies on supervised fine-tuning over datasets generated by LLMs.
Reinforcement Learning for Language Models. Reinforcement Learning (RL) (Kaelbling et al., 1996) is a machine learning paradigm that trains agents to make decisions through interaction with an environment, with the goal of maximizing cumulative rewards. In the era of large language models (LLMs), RL has gained significant traction as an effective framework for aligning models with human preferences. A prominent example is Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Stiennon et al., 2020; Ouyang et al., 2022), which leverages the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) and human preference data to train a reward model that guides the fine-tuning of LLMs. Building on RLHF, more recent algorithms such as GRPO (Shao et al., 2024) and REINFORCE++ (Hu, 2025) aim to enhance reward modeling and mitigate issues like biased optimization (Xu et al., 2024).
Reasoning with Language Models. It has been observed that sufficiently large language models (LMs) can demonstrate emergent reasoning capabilities (Wei et al., 2022; Suzgun et al., 2022). Chain-of-thought prompting (Wei et al., 2023) is one technique used to elicit step-by-step reasoning, significantly improving performance on complex tasks. Further advances include self-consistency (Wang et al., 2023) and structuring the reasoning process in forms like trees (Yao et al., 2023) or graphs (Besta et al., 2024; Cao, 2024) are also useful for more complex reasoning tasks. RL has also been used to directly improve reasoning skills during training (Lightman et al., 2023; Uesato et al., 2022). Notably, DeepSeek-R1 (DeepSeek-AI et al., 2025) demonstrates that large-scale RL can substantially boost the reasoning abilities of LMs. In terms of application, DeepRetrieval (Jiang et al., 2025) applies RL to teach models how to reason about interacting with search engines for information retrieval, while DeepCoder (Luo et al., 2025) uses RL for code reasoning and generation tasks. Rec-R1 (Lin et al., 2025) also bridges LLMs and recommendation systems through RL.
@@ 3 Methodology
In this section, we introduce our proposed training framework, Formula Tuning, illustrated in Figure 1. A summary of all notations is provided in Appendix M.
@@ 3.1 Task Formulation
where each $C_{i,j}$ may contain a data value, structural information (e.g., a top header or a left header), or be empty. In practice, we linearize $\mathbb{T}$ into a text sequence before feeding it to the LM. The LM then generates an output $a\in\mathcal{A}$, which can be either a final textual answer or a spreadsheet formula $f$ that produces the answer upon execution. Our goal is to optimize the parameters $\theta$ that maximize the expected table‐understanding performance, measured by a reward function $r(a\mid s)$. Formally,